# Privacy as Social Norm: Systematically Reducing Dysfunctional Privacy Concerns on Social Media

### v3

@@ Abstract.
Through co-design interviews ($N=19$) and a design evaluation survey ($N=136$) with U.S. teens ages 13-18, we investigated teens’ privacy management on social media. Our study revealed that 28% of teens with public accounts and 15% with private accounts experience dysfunctional fear, that is, fear that diminishes their quality of life or paralyzes them from taking necessary precautions. These fears fall into three categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps. While current approaches often emphasize individual vigilance and restrictive measures, our findings show this can paradoxically lead teens to either withdraw from beneficial social interactions or resign themselves to accept privacy violations, viewing them as inevitable. Drawing on teen input, we developed and evaluated ten design prototypes that emphasize empowerment over fear, system-wide explicit emphasis on privacy, clear privacy norms, and flexible controls. Survey results indicate teens perceive these approaches as effectively reducing privacy concerns while preserving social benefits. Our findings suggest that platforms will be more likely to protect teens’ privacy and less likely to manufacture unnecessary fear if they include designs that minimize the impact on other users, have low trade-offs with existing features, require minimal user effort, and function independently of community behavior. Such designs include: 1) alerting users about potentially unintentional personal information disclosure and 2) following up on user reports.
@@
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
In this paper, given the critical need for teen privacy protection in social media while preventing amplified dysfunctional fear, we investigate the following questions:
@@ 2.3. Toward a Community-level Privacy Support for Teens on Social Media
Echoing themes from fear of crime research, these phenomena underscore the need to understand worries and fears around privacy-related issues on social media. For example, fear of crime research differentiates between dysfunctional worry, which degrades the quality of life, and functional worry, which fosters precaution and awareness. Fear of crime research has found that only a quarter of worried individuals used their fear constructively, without impacting their life quality (Jackson and Gray, 2010). Fear of cybercrime studies reveal a similar pattern: while fear can be a motivator for protective behaviors, it often restricts activities and thus the quality of life, particularly when the fear response is disproportionate to the actual risk (Brands and Van Doorn, 2022) [^1] It is important to note, however, that our research does not aim to quantify the actual risks of privacy issues for teens on social media or determine what constitutes dysfunctional fear in this context. Rather, we seek to understand the emotional and psychological mechanisms of (dysfunctional) fear of crime and how they might apply to social media privacy concerns.. The influence of social environment on fear, as highlighted by the fear of crime research, may also be relevant to social media privacy concerns, suggesting that factors like (perceived) community disorder and (perceived) lack of social support can amplify fear (Lee et al., 2019; Greenberg, 1986; Lorenc et al., 2013). Therefore, in this paper, we argue that addressing social media privacy concerns requires a balance between vigilance and system-wide reassurance (Redmiles et al., 2019), as well as measures to increase self-efficacy. Following the “Privacy by Design” principle (Cavoukian, 2009), proactive approaches that integrate privacy into the design are essential. This is particularly important for teens, who are highly susceptible to peer pressure and social norms. However, it is crucial that community-wide privacy features do not merely create a “safety theater,” (Redmiles et al., 2019) which focuses solely on reducing fear without real safety, as a false sense of security can be dangerous (Cho et al., 2010; Walther, 2011; Acquisti and Gross, 2006; Gross and Acquisti, 2005; Ostendorf et al., 2020). The aim should be to empower users to move away from dysfunctional fear by transforming their concerns into constructive actions (Jackson and Gray, 2010).
@@ 4.1.4. Vague and Persistent Fear
Figure 1 of ten screenshots labeled from (a) to (j), each depicting different user interface features from a social media application focused on privacy: (a) Account protection: A settings menu allowing users to control who can follow them and comment on their posts, with options to limit interactions based on a "red flag" system. (b) Categorized viewer list: A feed section showing a post with a list of viewers categorized by the frequency of interaction. (c) On-demand screenshot blocking: A notification alerting the user that screenshots cannot be taken of a particular post because the poster has disabled this function. (d) Auto-delete (with save): A message interface with an option for the user to set their chats to auto-delete after a certain period while still allowing them to save specific messages. (e) Reminder to review: A feature offering users the option to decide the duration of their post’s visibility, ranging from 24 hours to permanent, with reminders to review this setting. (f) Red flag & follow-up: A notifications tab showing a red flag alert indicating that a post has been flagged and found to be in violation of community norms, with an option for more details. (g) Pseudonymity mode: A settings menu where users can activate a mode that hides their username and phone number on their public profile. (h) Personal information alert: An alert overlay on a photo within a post, warning the user that there appears to be a car plate number visible, and advising them to ensure no personal information is exposed. (i) User privacy norms: A community guidelines page highlighting the importance of respecting privacy and offering tips for safe engagement on the platform, with specific reference to the respect of personal space and keeping user information private. (j) "View as": A profile customization feature that allows users to preview their profile as it appears to other users, with options to see how it looks to users they follow and those they don’t, ensuring control over their public persona. Each screenshot is designed with a user-friendly interface, displaying clear text and interactive elements such as switches, dropdown menus, and informative pop-ups. These features collectively aim to enhance user privacy and provide personalized control over content sharing and interaction within the social media application.
@@
Enhanced security Pseudonymity mode (Fig. 1(g)) “Completely disconnects your real identity from your account, making it untraceable to certain people you choose, for ultimate privacy.”
@@ 4.2. How Does Design Affect Teens’ Privacy-related Fears?
On-demand screenshot blocking. This feature (Fig. 1(d)) addresses teens’ complex relationship with screenshot control. Our participants expressed nuanced views on screenshots, informed by their experiences with existing apps like Snapchat that notify users when screenshots are taken. They acknowledged that screenshots can threaten privacy by enabling out-of-context sharing, but also recognized legitimate reasons for taking them, such as preserving positive content. Participants noted that accidental screenshots can lead to awkward situations due to automatic notifications, and some preferred not to know when screenshots are taken to avoid overthinking. Given this range of preferences, we designed a feature allowing users to disable screenshots for individual posts. This approach aims to provide flexible control over content sharing, shape platform norms by conveying that screenshots are not always acceptable, and respect the original poster’s intentions.
@@ 6. Conclusion
In this paper, we explored the privacy-related fears experienced by teens on social media and investigated how design can mitigate these fears without compromising privacy protection. Our study involved co-design interviews with 19 teens and a design evaluation survey with 136 U.S. teen participants. Our findings indicate that teens frequently experience vague and persistent fears about privacy that they feel unable to manage effectively, leading to dysfunctional fear that inhibits their use of social media and negatively impacts their well-being. These fears fall into three categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps.
By incorporating system-wide guidance, enforcement, and options to make privacy both the norm and a priority, social media platforms can help teens navigate complex social interactions while maintaining privacy. Our study provides empirical evidence of dysfunctional fear in teens’ privacy management on social media and offers guidelines to support user empowerment. This work highlights the critical need to move beyond simply warning teens about privacy risks—which can lead to counterproductive anxiety and withdrawal—toward creating environments where privacy-conscious behavior is normalized and supported through thoughtful design.
Our findings emphasize that effective privacy protection for teens requires balancing safety with social connection, implementing proactive rather than fear-based approaches, and fostering community norms that support privacy-conscious behavior. Future research should continue to explore how platform design can help transform vague privacy anxieties into constructive privacy practices, enabling teens to fully participate in online social spaces while maintaining appropriate boundaries.
@@ Appendix A Demographics Data
Table 2. provides demographic information about the participants involved in the co-design study. It includes gender distribution, average age, and racial background of the participants. It describes the platforms that participants share content on frequently and comfortably. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (63%), Boys (26%), Non-binary or third gender (5%), Boy and non-binary or third gender (5%) Age 13 (10.5%), 14 (10.5%), 15 (15.8%), 16 (21.1%), 17 (21.1%), 18 (21.1%) Race White (53%), Asian or Asian-American (32%), Black or African American (11%), White and Asian or Asian-American (5%) Hispanic or Latin-American Origin No (84%), Yes (16%) Social Media Usage Instagram (100%), BeReal (84%), Snapchat (74%), Twitter (68%), TikTok (63%), Discord (11%), Reddit (11%), Tumblr (5%) Platforms Where Participants Reported Sharing Most Frequently Instagram (37%), BeReal (32%), Snapchat (11%)), Twitter (5%), YouTube (5%), Discord (5%), Rinsta (5%) Platforms Where Participants Reported Sharing Most Comfortably Instagram (26%), BeReal (26%), Snapchat (16%), Discord (16%), Rinsta (5%), Pinterest (5%), Finsta (5%)
@@
Table 3. contains information for the participants in the Follow-up (Design Evaluation) Survey. The table provides insights into the gender identity, mean age, race, Hispanic or Latin-American origin, and social media usage of the participants. These demographic details are essential for understanding the composition of the survey participants and are crucial for interpreting the survey results. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (59.6%), Boys (33.8%), Non-binary or third gender (2.9%), Girl and non-binary or third gender (1.5%), Boy, girl, and non-binary or third gender (1.5%), Prefer Not to Disclose (0.7%) Age 13 (0%), 14 (5.1%), 15 (18.4%), 16 (18.4%), 17 (29.4%), 18 (28.7%) Race White (47.8%), Black or African American (25.7%), Asian or Asian-American (20.6%), White and Black or African American (1.5%), White and Asian or Asian-American (0.7%), White and American Indian or Alaska Native (1.5%), Black or African-American, American Indian or Alaska Native (1.5%), Asian or Asian-American, Black or African-American (0.7%), Other (0.7%) Hispanic or Latin-American Origin No (83.1%), Yes (16.9%) Social Media Usage Instagram (Real) (85.3%), Snapchat (55.9%), TikTok (65.4%), BeReal (38.2%), Instagram (Spam) (39.7%), Twitter (40.4%) Other (7.9%): Reddit (2.9%), Youtube (1.5%), Pinterest (0.7%), Locket (0.7%), Lemon8 (0.7%), Pintrest (0.7%), Whatsapp (0.7%), Zeeme (0.7%)

### v4

@@ Abstract.
Through co-design interviews ($N=19$) and a design evaluation survey ($N=136$) with U.S. teens ages 13-18, we investigated teens’ privacy management on social media. Our study revealed that 28.1% of teens with public accounts and 15.3% with private accounts experience dysfunctional fear, that is, fear that diminishes their quality of life or paralyzes them from taking necessary precautions. These fears fall into three categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps. While current approaches often emphasize individual vigilance and restrictive measures, our findings show this can paradoxically lead teens to either withdraw from beneficial social interactions or resign themselves to accept privacy violations, viewing them as inevitable. Drawing on teen input, we developed and evaluated ten design prototypes that emphasize empowerment over fear, system-wide explicit emphasis on privacy, clear privacy norms, and flexible controls. Survey results indicate teens perceive these approaches as effectively reducing privacy concerns while preserving social benefits. Our findings suggest that platforms will be more likely to protect teens’ privacy and less likely to manufacture unnecessary fear if they include designs that minimize the impact on other users, have low trade-offs with existing features, require minimal user effort, and function independently of community behavior. Such designs include: 1) alerting users about potentially unintentional personal information disclosure and 2) following up on user reports.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW151publicationmonth: 4price: 15.00doi: 10.1145/3711049submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
In this paper, given the critical need for teen privacy protection in social media, while preventing amplified dysfunctional fear, we investigate the following questions:
@@ 2.3. Toward a Community-level Privacy Support for Teens on Social Media
Echoing themes from fear of crime research, these phenomena underscore the need to understand worries and fears around privacy-related issues on social media. For example, fear of crime research differentiates between dysfunctional worry, which degrades the quality of life, and functional worry, which fosters precaution and awareness. Fear of crime research has found that only a quarter of worried individuals used their fear constructively, without impacting their life quality (Jackson and Gray, 2010). Fear of cybercrime studies reveal a similar pattern: while fear can be a motivator for protective behaviors, it often restricts activities and thus the quality of life, particularly when the fear response is disproportionate to the actual risk (Brands and Van Doorn, 2022) [^1] It is important to note, however, that our research does not aim to quantify the actual risks of privacy issues for teens on social media or determine what constitutes dysfunctional fear in this context. Rather, we seek to understand the emotional and psychological mechanisms of (dysfunctional) fear of crime and how they might apply to social media privacy concerns.. The influence of social environment on fear, as highlighted by the fear of crime research, may also be relevant to social media privacy concerns, suggesting that factors like (perceived) community disorder and (perceived) lack of social support can amplify fear (Lee et al., 2019; Greenberg, 1986; Lorenc et al., 2013). Therefore, in this paper, we argue that addressing social media privacy concerns requires a balance between vigilance and system-wide reassurance (Redmiles et al., 2019), as well as measures to increase self-efficacy. Following the “Privacy by Design” principle (Cavoukian, 2009), proactive approaches that integrate privacy into the design are essential. This is particularly important for teens, who are highly susceptible to peer pressure and social norms. However, it is crucial that community-wide privacy features do not merely create a “safety theater,” (Redmiles et al., 2019) which focuses solely on reducing fear without real safety, as a false sense of security can be dangerous (Cho et al., 2010; Walther, 2011; Acquisti and Gross, 2006; Gross and Acquisti, 2005; Ostendorf et al., 2020). The aim should be to empower users to move away from dysfunctional fear by transforming their concerns into constructive actions (Jackson and Gray, 2010; Chordia et al., 2024).
@@ 4.1.4. Vague and Persistent Fear
Figure 1 of ten screenshots labeled from (a) to (j), each depicting different user interface features from a social media application focused on privacy: (a) Account protection: A settings menu allowing users to control who can follow them and comment on their posts, with options to limit interactions based on a "red flag" system. (b) Categorized viewer list: A feed section showing a post with a list of viewers categorized by the frequency of interaction. (c) On-demand screenshot blocking: A notification alerting the user that screenshots cannot be taken of a particular post because the poster has disabled this function. (d) Auto-delete (with save): A message interface with an option for the user to set their chats to auto-delete after a certain period while still allowing them to save specific messages. (e) Reminder to review: A feature offering users the option to decide the duration of their post’s visibility, ranging from 24 hours to permanent, with reminders to review this setting. (f) Red flag & follow-up: A notifications tab showing a red flag alert indicating that a post has been flagged and found to be in violation of community norms, with an option for more details. (g) Pseudonymity mode: A settings menu where users can activate a mode that hides their username and phone number on their public profile. (h) Personal information alert: An alert overlay on a photo within a post, warning the user that there appears to be a car plate number visible and advising them to ensure no personal information is exposed. (i) User privacy norms: A community guidelines page highlighting the importance of respecting privacy and offering tips for safe engagement on the platform, with specific reference to the respect of personal space and keeping user information private. (j) "View as": A profile customization feature that allows users to preview their profile as it appears to other users, with options to see how it looks to users they follow and those they don’t, ensuring control over their public persona. Each screenshot is designed with a user-friendly interface, displaying clear text and interactive elements such as switches, dropdown menus, and informative pop-ups. These features collectively aim to enhance user privacy and provide personalized control over content sharing and interaction within the social media application.
@@
Enhanced security Pseudonymity mode (Fig. 1(g)) “Completely disconnect your real identity from your account, making it untraceable to certain people you choose, for ultimate privacy.”
@@ 4.2. How Does Design Affect Teens’ Privacy-related Fears?
On-demand screenshot blocking. This feature (Fig. 1(d)) addresses teens’ complex relationship with screenshot control. Our participants expressed nuanced views on screenshots, informed by their experiences with existing apps like Snapchat that notify users when screenshots are taken. They acknowledged that screenshots could threaten privacy by enabling out-of-context sharing but also recognized legitimate reasons for taking them, such as preserving positive content. Participants noted that accidental screenshots can lead to awkward situations due to automatic notifications, and some preferred not to know when screenshots are taken to avoid overthinking. Given this range of preferences, we designed a feature allowing users to disable screenshots for individual posts. This approach aims to provide flexible control over content sharing, shape platform norms by conveying that screenshots are not always acceptable, and respect the original poster’s intentions.
@@ 6. Conclusion
We conducted co-design interviews with 19 teens and a design evaluation survey with 136 U.S. teen participants. The study revealed that teens experience significant dysfunctional privacy fears on social media, with 28.1% of public account users and 15.3% of private account users reporting fears that diminish their quality of life without enabling constructive protective actions. These fears stem from a perceived lack of control over audience reach, hostile online environments, and uncertainty about privacy norms.
Our work suggests that platforms should take greater responsibility for establishing privacy as a social norm rather than focusing solely on individual vigilance and restrictive measures. The ten design prototypes we evaluated demonstrate concrete ways to empower users while reducing unnecessary anxiety. The features that were perceived as most effective shared common characteristics: They minimized impact on other users, had low trade-offs with existing functionality, required minimal user effort, and functioned independently of community behavior. Particularly successful examples included alerts about potential privacy exposures and systematic follow-up on user reports.
The implications of this work extend beyond specific features to suggest a broader shift in how platforms approach teen privacy. Rather than emphasizing individual responsibility and vigilance, social media platforms should work to establish privacy protection as a community norm through system-level design choices. This involves not only providing robust privacy controls but also creating an environment where privacy-conscious behavior is explicitly valued and supported. Future research should continue to explore how platform design can help transform vague privacy anxieties into constructive privacy practices, enabling teens to fully participate in online social spaces while maintaining appropriate boundaries.
@@ Appendix A Demographics Data
Table 2. provides demographic information about the participants involved in the co-design study. It includes gender distribution, average age, and racial background of the participants. It describes the platforms that participants share content on frequently and comfortably. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (63%), Boys (26%), Non-binary or third gender (5%), Boy and non-binary or third gender (5%) Age 13 (10.5%), 14 (10.5%), 15 (15.8%), 16 (21.1%), 17 (21.1%), 18 (21.1%) Race White (53%), Asian or Asian-American (32%), Black or African American (11%), White and Asian or Asian-American (5%) Hispanic or Latin-American Origin No (84%), Yes (16%) Social Media Usage Instagram (100%), BeReal (84%), Snapchat (74%), Twitter (68%), TikTok (63%), Discord (11%), Reddit (11%), Tumblr (5%) Platforms Where Participants Reported Sharing Most Frequently Instagram (42%), BeReal (32%), Snapchat (11%)), Twitter (5%), YouTube (5%), Discord (5%) Platforms Where Participants Reported Sharing Most Comfortably Instagram (36%), BeReal (26%), Snapchat (16%), Discord (16%), Pinterest (5%)
@@
Table 3. contains information for the participants in the Follow-up (Design Evaluation) Survey. The table provides insights into the gender identity, mean age, race, Hispanic or Latin-American origin, and social media usage of the participants. These demographic details are essential for understanding the composition of the survey participants and are crucial for interpreting the survey results. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (59.6%), Boys (33.8%), Non-binary or third gender (2.9%), Girl and non-binary or third gender (1.5%), Boy, girl, and non-binary or third gender (1.5%), Prefer Not to Disclose (0.7%) Age 13 (0%), 14 (5.1%), 15 (18.4%), 16 (18.4%), 17 (29.4%), 18 (28.7%) Race White (47.8%), Black or African American (25.7%), Asian or Asian-American (20.6%), White and Black or African American (1.5%), White and Asian or Asian-American (0.7%), White and American Indian or Alaska Native (1.5%), Black or African-American, American Indian or Alaska Native (1.5%), Asian or Asian-American, Black or African-American (0.7%), Other (0.7%) Hispanic or Latin-American Origin No (83.1%), Yes (16.9%) Social Media Usage Instagram (Real) (85.3%), Snapchat (55.9%), TikTok (65.4%), BeReal (38.2%), Instagram (Spam) (39.7%), Twitter (40.4%) Other (7.9%): Reddit (2.9%), Youtube (1.5%), Pinterest (0.7%), Locket (0.7%), Lemon8 (0.7%), Whatsapp (0.7%), Zeeme (0.7%)